OpenStudy (anonymous):

anyone want to help? Algebra 2 questions!!

OpenStudy (anonymous):

I'll have a crack

OpenStudy (anonymous):

it's in the attachment

OpenStudy (anonymous):

Well I can't really remember the trigonometric identities for question 1, but question 2, the sequence looks pretty simple. The top row increases by 1 while bottom is the odd numbers.\[\frac{ n }{ 2n-1 }\]Is the sequence. Question 3 is halving and changing sign on each iteration.\[\frac{ (-1)^{n+1}16 }{ 2^n }\]I think it's something like that (it's worth double checking. For the graph, the amplitude is simply the extreme (in this case I think it is 4?) and the period is how long it take to repeat an entire wave (in this case I think it is \(\pi\). Hope that helps, let me know if you have any problems!

OpenStudy (anonymous):

yeah thanks!